Grade 12th PassPhysical Chemistry

If the incident frequency of a radiation is doubled,then the stopping potential will be :
A)Halved
B)Doubled
C)More than doubled
D)Less than doubled

Manas Mishra
8 Years agoGrade 12th Pass

1 Answer

Khimraj
8 Years ago
The photoelectric eqn is: 

hf - φ = qVs, 

where, the stopping potential can be solved; 

Vs = (hf - φ)/q 

let f' = 2f 

and work function: φ of the metal is constant. 

∆Vs = Vs' - Vs = [(h*2f - φ) - (hf - φ)]/q = hf/q 

while, 

2Vs = 2(hf - φ))/q 

so that, 

∆Vs > 2Vs 

meaning that if the frequency is doubled the 
stopping potential, Vs increases to more than doubled.